Calculus

1.1 Find the area of the region that lies inside region that lies inside r= cos 20 and outside r=Find the volume of the parallelepiped determined by a=< 1, 2, -1>, b=-2i+3 k and c=73–4k.2.
Answer #1

1. ra cos(20) and =% Now, Los (20) =Y₂ % r = cos(28 (20) =/ Now, area of the region that lies inside r = cos(20) and outsidea=(1,2,-1) b = -zitsk Then an (o xe) ca 7 -4K keta bo 1 2 - 1 ܙܙ -2 O 3 07-4 ! (0-212-2(8-0) -1(-14) - 21-16 +14 17 - 23 volu